# CNIMA: A Universal Evaluation Framework and Automated Approach for Assessing Second Language Dialogues

### v1

@@ Abstract
We develop CNIMA (Chinese Non-Native Interactivity Measurement and Automation), a Chinese-as-a-second-language labelled dataset with 10K dialogues. We annotate CNIMA using an evaluation framework — originally introduced for English-as-a-second-language dialogues — that assesses micro-level features (e.g. backchannels) and macro-level interactivity labels (e.g. topic management) and test the framework’s transferability from English to Chinese. We found the framework robust across languages and revealed universal and language-specific relationships between micro-level and macro-level features. Next, we propose an approach to automate the evaluation and find strong performance, creating a new tool for automated second language assessment. Our system can be adapted to other languages easily as it uses large language models and as such does not require large-scale annotated training data.
CNIMA: A Universal Evaluation Framework and
Automated Approach for Assessing Second Language Dialogues
@@ 1 Introduction
In the context of second language (SL) assessment, speaking has always been considered an essential ability for SL speakers Smith et al. (2022); Allwood (2008); Huang et al. (2020), but prior studies have predominantly focused on written language proficiency Paiva et al. (2022). Given the lack of datasets that capture the unique linguistic features of SL speakers in dialogues (especially in open-domain conversations), this leaves conversational interaction assessment under-explored and ultimately contribute to a limited understanding of conversational fluency and interactivity for SL speakers.
@@
An exception is Gao et al. (2024), who propose a two-level framework for assessing the interactivity ability of English-as-a-second-language (ESL) speakers in open-domain conversations. They introduce micro-level word/utterance features (e.g. backchannels) and macro-level interactivity labels (e.g. topic management) and found that the micro-level features are highly predictive of the macro-level labels. Gao et al. (2024) reveal their intricate relationships, showing that certain micro-level features are highly correlated with a particular macro-level feature, e.g. frequent use of appropriate collocations and varied sentence structures (micro-level features) significantly correlate with topic management and tone appropriateness (macro-level labels). However, this research is conducted only in the context of ESL, raising questions about the transferability of the evaluation framework to other languages besides English. Furthermore, they did not propose an automated pipeline for assessing SL dialogue quality, as their predictive models rely on human-annotated micro-level features.
Our work aims to address these shortcomings by (1) developing and testing a Chinese-as-a-second-language (CSL) dialogue dataset using the evaluation framework of Gao et al. (2024); (2) introducing a fully automated pipeline to assess interactivity for CSL conversations. We experiment with different machine learning models, from classical machine learning models (e.g. logistic regression) to large language models (LLMs; e.g. GPT-4o OpenAI (2024)), to automate the prediction of micro/macro-level features and overall dialogue quality scores. Our best system shows strong predictive performance, paving the way for a new tool for assessing SL conversations [^1] The dataset and related codes can be accessed at https://github.com/RenaGao/CSL2024. To summarise:
@@
  • We release CNIMA (Chinese Non-Native Interactivity Measurement and Automation), an annotated CSL dialogue dataset with 10K dialogues, based on the framework of Gao et al. (2024) originally designed for ESL.
@@
  • We test the cross-lingual transferability of the evaluation framework for CSL and find that it is robust across languages, and we further reveal universal and language-specific relationships between the micro-level and macro-level features.
@@
  • We introduce a fully automated approach that predicts micro-level features, macro-level labels and an overall quality score given an input dialogue by exploring a spectrum of models from classical machine learning to LLMs. Our best system demonstrates strong performance, creating a new automatic tool for second language assessment. Importantly, our LLM-based system can be adapted to other languages without requiring annotated data.
@@ 2.1 Assessment of SL Spoken Conversation
Mainstream second language assessments in current industrial practice have mainly focused on grammatical accuracy, pronunciation standardization and vocabulary richness; for example, in TOEFL iBT, PTE Academic and Cambridge IELTS test Xu (2018); Paiva et al. (2022); Xu et al. (2021). However, few speaking assessments emphasised the importance of interaction Khabbazbashi et al. (2021) and aspects such as how speakers manage the topics in communication Shaxobiddin (2024), how speakers perform social roles from speaking Chen et al. (2023), and how speakers start and end a talk in an acceptable manner Yap and Sahoo (2024). There are, however, some exceptions. For example, Dai (2022) develop a test rubric for Chinese SL speakers. More recently, Gao and Wang (2024) introduce an interactivity scoring framework inspired by the IELTS speaking assessment. However, these studies only provide a theoretical framework for evaluation and lack an automated pipeline for large-scale assessments.
@@ 2.2 Automated Scoring System in SL
Automated scoring systems can improve the efficiency of processing and scoring dialogues compared to manual assessment methods, allowing for real-time feedback Evanini et al. (2017). Some major second language assessment organisations have employed automated speaking scoring, including PTE Pearson Jones and Liu (2023), Duolingo English Test Burstein et al. (2021), and TOEFL Gong (2023), but these automated systems fail to capture the interactive nature of SL conversations. That is, they struggle to offer detailed analyses and insights on the common errors and language usage patterns and where SL speakers struggle the most. This motivates a more explainable framework for SL assessment that can provide better feedback.
The main challenge is in designing an effective evaluation framework. Capturing the nuances of spoken language, such as interaction, attitudes, and cooperation in conversation, is complex to achieve Gao and Wang (2024). Automated scoring systems must handle diverse patterns, which can vary widely among different non-native SL speakers. Recognizing and assessing interactive features like turn-taking, interruptions, and response appropriateness further complicates the process. Ultimately, the dynamic nature of conversations makes it challenging for automated systems to accurately evaluate interactive speaking, necessitating sophisticated algorithms and continuous refinement Cumbal (2024); Engwall et al. (2022).
@@ 3 Evaluation Framework
We adapt the dialogue evaluation framework of Gao et al. Gao et al. (2024) for assessing SL dialogues, which considers micro-level features and macro-level interactivity labels. One addition we have made is to introduce an overall dialogue quality score. Figure 1 shows an example of CSL dialogue annotation with the evaluation framework.
The framework has 17 micro-level features, and it can be further broken into token-level (such as ‘Reference Word’: she and ‘Backchannel’: hmm) and utterance-level features (such as ‘Formulaic response’: How’s going). The micro-level features are annotated as spans (i.e. annotators mark text spans corresponding to a feature). For macro-level interactivity labels, there are four: Topic Management, Tone Choice Appropriateness, Conversation Opening and Conversation Closing, and they are annotated as (dialogue) labels. Briefly, topic management refers to the speaker’s ability to control and navigate the flow of topics; tone choice the suitability of the tone; and conversation opening/closing the naturalness in the initial exchange and conclusion of the discussion. A more elaborate definition can be found in Appendix Table 13. For each macro-level interactivity label, the score ranges from 1 to 5 (categorical), and higher scores indicate a more natural and authentic interactivity quality. Note that for tone choice appropriateness, higher scores denote a casual tone, and lower scores indicate a formal tone. [^2] The rationale of assigning higher scores to casual tone is that in our experiments, the conversations are designed to be informal discussions, and appropriate use of casual tone signifies a more natural communicative interaction.
In addition to micro- and macro-level features, we introduce an overall score for each dialogue to measure second language speakers’ interaction abilities in open-domain conversation. Like the macro-level labels, it is scored from 1 to 5 (categorical). The overall score is designed to capture the holistic quality of a conversation, integrating elements like contextualisation, responsiveness, and communicative purposes across the whole conversation, and a high score reflects the speaker’s ability to engage in fluid and meaningful interactions. The full description of each score is detailed in Appendix Table 12.
@@ 4.1 CSL Dialogue Collection
We first extend the CSL dialogue dataset developed by Wu and Roever (2021), which has approximately 8,000 dialogues. We followed a similar process, where we recruited 20 learners of Chinese Mandarin at four different proficiency levels. [^3] Proficiency level can be categorized into the following levels: beginner, lower, intermediate, and high Wu and Roever (2021). Each participant was assigned to a dyadic group with a partner of a similar proficiency level. Each pair of participants did one elicited conversation task, in which they discussed an assigned topic, and two role-play activities (three tasks in total; dialogue collection instructions can be found in Appendix LABEL:subapp:SI). [^4] All participants have signed consent approved by an ethics board to agree to audio recording. After collecting the conversations, we recruited in-house workers to segment the conversations based on the discussion topics. [^5] The workers are undergraduate linguistics major students, and they are native Chinese speakers. Including the original 8,000 dialogues from Wu and Roever (2021), our extended CSL dataset has 10,908 dialogues in total; table 1 presents some statistics.
@@
4.2 Macro/Micro-level Feature Annotation
@@
Given the dialogues, we annotate them for micro-level features, macro-level interactivity labels and overall quality scores based on the evaluation framework introduced in Section 3. To this end, we have recruited twelve postgraduate students who are native Chinese speakers. The authors of this paper first trained the annotators and the annotator training manual can be found in Appendix  LABEL:subapp:MN. During the training, annotators will see one dialogue as an example to understand the requirements and learn how to use the annotation platform (Appendix A.1). After training, each annotator was assigned 950 dialogues, and each dialogue was annotated by two annotators. [^6] The assignment is created in a way such that we have the same pair of annotators for every batch of 40 dialogues. The annotation was conducted over two months, where in the first few weeks we (first author) checked for initial agreement, discussed feedback, and fixed any annotation errors (e.g. missing values in the annotation results) before proceeding with the full annotation. The annotation process was guided by an annotation guide (Appendix LABEL:subapp:MN), which provided definitions and examples for each micro-level feature, macro-level interactivity label and the overall dialogue quality score. For the overall score, in addition to providing the score, the annotators are also asked to write down their reasons for justifying their label.
@@
As explained in Section 3, the micro-level features are annotated as spans, and the macro-level labels and overall quality scores are document (dialogue) labels. To aggregate the span annotations by the two annotators for the micro-level features, for each dialogue and each micro-level feature, we iterate through each turn and select the shorter (longer) span between the two annotators if it is a token-level (utterance-level) feature. [^7] We do this because token-level features (e.g., ‘Reference word’ (She, her, he) or ‘Tense Choice’ (is doing, done, did) tends to be very short and limited to 1 or 2 words and so the shorter spans are likely to be more accurate Greer (2023). The same reasoning applies to utterance-level spans. For the macro-level interactivity labels, in dialogues where we have disagreement between the two annotators, we use the majority label from their larger unsegmented conversation as the ground truth label. [^8] Recall that these dialogues are segmented from a larger, longer conversation. The majority label is the most frequent label when we pool together all the labels from the segmented, smaller dialogues that belong to the same original conversation. Our rationale for deferring to this majority label is that these are cases where it is difficult to determine the correct dialogue level label using the smaller segmented dialogue (this happens more with topic management than other interactivity labels), and so it is best that we look at all the labels across all the segmented dialogues collection that belong to the same conversation. For the overall dialogue quality scores, for each dialogue with disagreement, we (two authors of this paper) manually assess the justification provided by the annotators to determine the ground truth. [^9] Note that when resolving the disagreement, we only consider the justification without looking at their macro-level interactivity labels. This is so we create an unbiased ground truth for the overall score that is independent of the macro-level labels.
@@
To measure annotation quality, for the micro-level features, we measure agreement between the annotators at the token level for each micro-level feature, i.e., we compute agreement statistics based on the presence or absence of the feature as marked by the annotators for each word token. [^10] In other words, the unit of analysis here is a word token, and the output is a binary value for each annotator indicating whether it has been marked for the feature. We calculate Pearson correlation coefficient $r$ Cohen et al. (2009) and Krippendorff’s $\alpha$ Krippendorff (2018) summarise the results in Table 2. The agreement is above 0.6 over all the features/labels, indicating a good consensus between annotators under this framework.
@@ 5 Automation Pipeline
This step predicts the macro-level interactivity labels for topic management, tone choice appropriateness, and conversation opening and closing. Here, we experiment with classical models (LR, RF and NB), BERT and GPT-4o, and for each of them, we train four (one for each interactivity label) 5-class classifiers (each interactivity label has 5 classes). For the classical models, we follow Gao et al. (2024) and convert the micro-level feature spans into normalised counts and use them as input features to train LR/RF/NB to predict the interactivity labels. For BERT and GPT-4o, the input is the dialogue concatenated with the micro-level feature spans. BERT is fine-tuned as a document classifier, and GPT-4o is one-shot prompted with instructions.
@@
This step predicts the overall dialogue quality score. As this is a document classification like step 2, we follow largely the same process for building LR, RF, NB, BERT and GPT-4o. Note that the classical models (LR, RF, NB) use only the normalised macro-level interactivity score as input features (in other words, they have only 4 input features), while BERT and GPT-4o uses a concatenation of the dialogue, micro-level feature spans, and macro-level interactivity scores as input.

### v2

# An Interpretable and Crosslingual Method for Evaluating Second-Language Dialogues

@@ Abstract
We analyse the cross-lingual transferability of a dialogue evaluation framework that assesses the relationships between micro-level linguistic features (e.g. backchannels) and macro-level interactivity labels (e.g. topic management), originally designed for English-as-a-second-language dialogues. To this end, we develop CNIMA (Chinese Non-Native Interactivity Measurement and Automation), a Chinese-as-a-second-language labelled dataset with 10K dialogues. We found the evaluation framework to be robust across distinct languages: English and Chinese, revealing language-specific and language-universal relationships between micro-level and macro-level features. Next, we propose an automated, interpretable approach with low data requirement that scores the overall quality of a second-language dialogue based on the framework. Our approach is interpretable in that it reveals the key linguistic and interactivity features that contributed to the overall quality score. As our approach does not require labelled data, it can also be adapted to other languages for second-language dialogue evaluation.
An Interpretable and Crosslingual Method for Evaluating Second-Language Dialogues
@@ 1 Introduction
In the context of second language (SL) assessment, speaking has always been considered an essential ability Smith et al. (2022); Allwood (2008); Huang et al. (2020), but prior studies have predominantly focused on written correction Paiva et al. (2022) or pronunciation from ASR McGhee et al. (2024). Given the lack of researches that capture the unique linguistic features of SL speakers in dialogues (especially in open-domain interactive conversations), this leaves conversational interaction assessment under-explored and ultimately contributes to a limited understanding of conversational fluency and interactivity for SL speakers from different languages.
@@
An exception is Gao et al. (2025), who propose a two-level framework for assessing the interactivity ability of English-as-a-second-language (ESL) speakers in open-domain conversations. The framework introduces micro-level word/utterance features (e.g. backchannels) and macro-level interactivity labels (e.g. topic management) and they find that the micro-level features are highly predictive of the macro-level labels. However, they conduct the study only in the context of ESL, raising questions about the transferability of the evaluation framework to other languages besides English. Furthermore, they do not introduce an automated approach for this evaluation, as the analysis relies on human annotations.
Our work aims to address these shortcomings by: (1) testing the cross-lingual transferability of the evaluation framework by validating it on an annotated Chinese-as-a-second-language (CSL) dialogue dataset; and (2) introducing an automated, interpretable approach with low data requirement that scores the overall quality of a second-language dialogue. Our automated approach is interpretable as it highlights the key features that lead to the overall score, and it has a low data requirement because it does not require labelled data. As it can be adapted to other languages easily, our method paves the way for a new tool for assessing SL conversations. [^1] The dataset and related codes can be accessed at https://github.com/RenaGao/CSL2024 To summarise our contributions:
@@
  • We release CNIMA (Chinese Non-Native Interactivity Measurement and Automation), an annotated CSL dialogue dataset with 10K dialogues, based on the evaluation framework of Gao et al. (2025) originally designed for ESL dialogues.
@@
  • We evaluate the cross-lingual transferability of the evaluation framework and find it to be robust across different languages. We further reveal language-specific and language-universal relationships between micro-level and macro-level features, highlighting the subtle differences between CSL and ESL dialogues.
@@
  • We introduce an automated, interpretable approach that predicts the micro- and macro-level features and the overall quality scores of second-language dialogues. Our best method demonstrates strong performance, creating a new tool for second language assessment. Importantly, the method can be adapted to other languages as it does not require labelled data.
@@ 2.1 Assessment of SL Spoken Conversation
Mainstream second language assessments in current industrial practice have mainly focused on grammatical accuracy, pronunciation standardization and vocabulary richness; for example, in TOEFL iBT, PTE Academic and Cambridge IELTS test Xu (2018); Paiva et al. (2022); Xu et al. (2021). However, few speaking assessments emphasised the importance of interaction in dialogues for second language speakers and learners Khabbazbashi et al. (2021) and aspects such as how speakers manage the topics in communication Shaxobiddin (2024), how speakers perform social roles from speaking interactions Chen et al. (2023), and how speakers start and end a talk in an acceptable manner Yap and Sahoo (2024). There are, however, some exceptions. For example, Dai (2022) develop a test rubric for Chinese Second Language speakers. More recently, Gao and Wang (2024) introduce an interactivity scoring framework inspired by the IELTS speaking assessment. However, these studies only provide a theoretical framework for evaluation and lack an automated pipeline for large-scale assessments, human rating is still a major scoring approach in mainstream industrial practice.
@@ 2.2 Automated Scoring System in SL
Automated scoring systems can improve the efficiency of processing and scoring dialogues compared to manual assessment methods, allowing for real-time feedback Evanini et al. (2017) and explainable suggestions. Some major second language assessment organisations have employed automated speaking scoring, including PTE Pearson Jones and Liu (2023), Duolingo English Test Burstein et al. (2021), and TOEFL Gong (2023), but these automated systems failed to capture the interactive nature of SL conversations. That is, they struggle to offer detailed analyses and insights on the common errors and language usage patterns and where SL speakers struggle the most. This motivates a more explainable framework for SL assessment that can provide better feedback and suggestions from the automated scoring.
The main challenge is to design an effective automated evaluation tool. Capturing the nuances of spoken language, such as interaction, attitudes, and cooperation in conversation, is challenging to achieve Gao and Wang (2024). Automated scoring systems must handle diverse patterns across varied languages, which can vary widely among different non-native SL speakers across distinct languages, like Asian languages and European languages with distinct differences. Assessing interactive features like turn-taking, interruptions, and response appropriateness further complicates the process. Ultimately, the dynamic nature of conversations makes it challenging for automated systems to accurately evaluate interactive speaking, necessitating sophisticated algorithms and continuous refinement Engwall et al. (2022); Cumbal (2024).
@@ 3 Evaluation Framework
We extend the evaluation framework of Gao et al. (2025) for assessing second-language dialogues, which considers micro-level features and macro-level interactivity labels. One addition we have made is to introduce an overall dialogue quality score. Figure 1 shows an example of CSL dialogue annotation with the evaluation framework. The framework has 17 micro-level features, and it can be further broken into token-level (such as ‘Reference Word’: she and ‘Backchannel’: hmm) and utterance-level features (such as ‘Formulaic response’: How’s going). The micro-level features are annotated as spans (i.e. annotators mark text spans corresponding to a feature). For macro-level interactivity labels, there are four: Topic Management, Tone Choice Appropriateness, Conversation Opening and Conversation Closing, and they are annotated as (dialogue) labels. Briefly, Topic Management refers to the speaker’s ability to control and navigate the flow of topics; Tone Choice the suitability of the tone; and Conversation Opening/Closing the naturalness in the initial exchange and conclusion of the discussion. A more elaborate definition can be found in Appendix Table 13. For each macro-level interactivity label, the score ranges from 1 to 5 (categorical), and higher scores indicate a more natural and active interactivity quality. Note that for tone choice appropriateness, higher scores denote a casual tone, and lower scores indicate a formal tone. [^2] The rationale of assigning higher scores to casual tone is that in our experiments, the conversations are designed to be informal discussions, and appropriate use of casual tone signifies a more natural communicative interaction. In addition to micro- and macro-level features, we introduce an overall score for each dialogue to measure second language speakers’ interaction abilities in open-domain conversation. Like the macro-level labels, it is scored from 1 to 5 (categorical). The overall score is designed to capture the holistic quality of a conversation, integrating elements like contextualisation, responsiveness, and communicative purposes across the whole conversation, and a high score reflects the speaker’s ability to engage in fluid and meaningful interactions. The full description of each score is detailed in Appendix Table 12.
@@ 4.1 CSL Dialogue Collection
We first extend the CSL dialogue dataset developed by Wu and Roever (2021), which has approximately 8,000 dialogues. We followed a similar process, where we recruited 20 learners of Chinese Mandarin at four different proficiency levels. [^3] Proficiency level can be categorized into beginner, lower, intermediate, and high Wu and Roever (2021). Each participant was assigned to a dyadic group with a partner of a similar proficiency level. Each pair of participants did one elicited conversation task, in which they discussed an assigned topic, and two role-play activities (three tasks in total; dialogue collection instructions can be found in Appendix LABEL:subapp:SI). [^4] All participants have signed consent approved by an ethics board to agree to audio recording. After collecting the conversations, we recruited in-house workers to segment the conversations based on the discussion topics. [^5] The workers are undergraduate linguistics major students, and they are native Chinese speakers. Including the original 8,000 dialogues from Wu and Roever (2021), our extended CSL dataset has 10,908 dialogues in total as shown in Table 1.
@@
4.2 Data Annotation
@@
Given the dialogues, we annotate them for micro-level features, macro-level interactivity labels and overall quality scores based on the evaluation framework introduced in Section 3. To this end, we recruited twelve postgraduate students who are native Chinese speakers. We (first author) first trained the annotators and the annotator training manual can be found in Appendix  LABEL:subapp:MN. During the training, annotators will see one dialogue as an example to understand the requirements and learn how to use the annotation platform (Appendix A.1). After training, each annotator was assigned 950 dialogues, and each dialogue was annotated by two annotators. [^6] The assignment is created in a way such that we have the same pair of annotators for every batch of 40 dialogues. The annotation was conducted over two months, where in the first few weeks we (first author) checked for initial agreement, discussed feedback, and fixed any annotation errors (e.g. missing values in the annotation results) before proceeding with the full annotation. The annotation process was guided by an annotation guide (Appendix LABEL:subapp:MN), which provided definitions and examples for each micro-level feature, macro-level interactivity label and the overall dialogue quality score. For the overall score, in addition to providing the score, the annotators are also asked to write down their reasons for justifying their label.
@@
As explained in Section 3, the micro-level features are annotated as spans, and the macro-level labels and overall quality scores as document (dialogue) labels. To aggregate the span annotations by the two annotators for the micro-level features, for each dialogue and each micro-level feature, we iterate through each turn and select the shorter (longer) span between the two annotators if it is a token-level (utterance-level) feature. [^7] We do this because token-level features (e.g., ‘Reference word’ (She, her, he) or ‘Tense Choice’ (is doing, done, did) tends to be very short and limited to 1 or 2 words and so the shorter spans are likely to be more accurate Greer (2023). The same reasoning applies to utterance-level spans. For the macro-level interactivity labels, in dialogues where we have disagreement between the two annotators, we use the majority label from their larger unsegmented conversation as the ground truth label. [^8] Recall that these dialogues are segmented from a larger, longer conversation. The majority label is the most frequent label when we pool together all the labels from the segmented, smaller dialogues that belong to the same original conversation. Our rationale for deferring to this majority label is that these are cases where it is difficult to determine the correct dialogue level label using the smaller segmented dialogue (this happens more with topic management than other interactivity labels), and so we look at all the labels across all the segmented dialogues collection that belong to the same conversation. For the overall dialogue quality scores, for each dialogue with disagreement, we (two authors of this paper) manually assess the justification provided by the annotators to determine the ground truth. [^9] Note that when resolving the disagreement, we only consider the justification without looking at their macro-level interactivity labels. This is so we create an unbiased ground truth for the overall score that is independent of the macro-level labels.
@@
To measure the annotation quality, for the micro-level features, we calculate agreement between the annotators at the token level for each micro-level feature, i.e., we compute agreement statistics based on the presence or absence of the feature as marked by the annotators for each word token. [^10] In other words, the unit of analysis here is a word token, and the output is a binary value for each annotator indicating whether it has been marked for the feature. We calculate Pearson correlation coefficient $r$ Cohen et al. (2009) and Krippendorff’s $\alpha$ Krippendorff (2018) and summarise the results in Table 2. The agreement is above 0.6 over all the features/labels, indicating a good consensus between annotators.
@@ 5 Automated Pipeline
This step predicts the macro-level interactivity labels for topic management, tone choice appropriateness, and conversation opening and closing. Here, we experiment with classical models (LR, RF and NB), BERT and GPT-4o. For each of them, we train four (one for each interactivity label) 5-class classifiers (each label has 5 classes). For the classical models, we follow Gao et al. (2025) and convert the micro-level feature spans into normalised counts and use them as input features to train LR/RF/NB to predict the interactivity labels. For BERT and GPT-4o, the input is the dialogue concatenated with the micro-level feature spans. BERT is fine-tuned as a document classifier, and GPT-4o is one-shot prompted with instructions.
@@
This step predicts the overall dialogue quality score. As this is a document classification task like step 2, we follow the same process for building LR, RF, NB, BERT and GPT-4o. Note that the classical models (LR, RF, NB) use only the normalised macro-level interactivity score as input features (in other words, they have only 4 input features), while BERT and GPT-4o uses a concatenation of the dialogue, micro-level feature spans, and macro-level interactivity scores as input.